\begin{thebibliography}{ELMNP}
%\bibitem[Art69]{Art69} M. Artin.  {\it Algebraization of formal moduli. I.} In Global Analysis (Papers in honor of K. Kodaira), pages 21–71. Univ. Tokyo Press, Tokyo, 1969.
\bibitem[BCHM10]{BCHM10} Caucher Birkar, Paolo Cascini, Christopher D. Hacon, and James McKernan.  {\it Existence of minimal models for varieties of log general type.} J. Amer. Math. Soc., 23(2):405–468, 2010. 
%\bibitem[BCZ15]{BCZ15} Caucher Birkar, Yifei Chen, and Lei Zhang.  {\it Iitaka’s $C_{n,m}$ conjecture for 3-folds over finite fields.} arXiv preprint arXiv:1507.08760, 2015.
\bibitem[Bir15]{Bir15} Caucher Birkar.  {\it Existence of flips and minimal models for 3-folds in char $p$.} arXiv preprint arXiv:1311.3098, 2015.
%\bibitem[BLR12]{BLR12} Siegfried Bosch, Werner Lutkebohmert, and Michel Raynaud.  {\it N\'eron models,} volume 21. Springer Science and Business Media, 2012.
%\bibitem[BP12]{BP12} Bo Berndtsson and Mihai P˘aun.  {\it Quantitative extensions of pluricanonical forms and closed positive currents.} Nagoya Math. J., 205:25–65, 2012.
%\bibitem[CL12]{CL12} Paolo Cascini and Vladimir Lazi\'c.  {\it New outlook on the minimal model program, I.} Duke Math. J., 161(12):2415–2467, 2012.
\bibitem[CL13]{CL13} Alessio Corti and Vladimir Lazi\'c.  {\it New outlook on the minimal model program, II.} Math. Ann., 356(2):617-–633, 2013.
%\bibitem[CL14]{CL14} Paolo Cascini and Vladimir Lazi\'c.  {\it On the number of minimal models of a log smooth threefold.} J. Math. Pures Appl. (9), 102(3):597–616, 2014.
%\bibitem[CP09]{CP09} Vincent Cossart and Olivier Piltant.  {\it Resolution of singularities of threefolds in positive characteristic. II.} J. Algebra, 321(7):1836–1976, 2009.
%\bibitem[CZ14]{CZ14} Paolo Cascini and De-Qi Zhang.  {\it Eﬀective ﬁnite generation for adjoint rings.} Ann. Inst. Fourier (Grenoble), 64(1):127–144, 2014.
%\bibitem[DCF15]{DCF15} Gabriele Di Cerbo and Andrea Fanelli.  {\it Eﬀective Matsusaka’s theorem for surfaces in characteristic $p$.} Algebra Number Theory, 9(6):1453–1475, 2015.
%\bibitem[DERU13]{DERU13} Pierre D\'ebes, Michel Emsalem, Matthieu Romagny, and A. Muhammed Uludag, editors.  {\it Arithmetic and geometry around Galois theory,} volume 304 of Progress in Mathematics. Birkh¨auser/Springer, Basel, 2013. Lecture notes from the summer schools Geometry and Arithmetic of Moduli Spaces of Covering held at Galatasaray University, Istanbul, June 9–20, 2008, and Geometry and Arithmetic around Galois Theory held at Galatasaray University, Istanbul, June 8–19, 2009.
%\bibitem[dFE14]{dFE14} Tommaso de Fernex and Lawrence Ein.  {\it Vanishing theorems and singularities in birational geometry.} 2014.
\bibitem[Egb16]{Egb16} Andrew Egbert. {\it Log Minimal Models For Arithmetic Threefolds}, {(2016)}, {\text url: {https://github.com/divbit/phd-thesis}}.
%\bibitem[Eke88]{Eke88} Torsten Ekedahl.  {\it Canonical models of surfaces of general type in positive characteristic.} Inst. Hautes \'Etudes Sci. Publ. Math., (67):97–144, 1988.
%\bibitem[ELM+06]{ELM+06} Lawrence Ein, Robert Lazarsfeld, Mircea Mustata, Michael Nakamaye, and Mihnea Popa.  {\it Asymptotic invariants of base loci.} Ann. Inst. Fourier (Grenoble), 56(6):1701–1734, 2006.
%\bibitem[GD71]{GD71} A. Grothendieck and J. A. Dieudonn\'e.  {\it El\'ements de g\'eom\'etrie alg\'ebrique.} I, volume 166 of Grundlehren der Mathematischen Wissenschaften [Fundamental Principles of Mathematical Sciences]. Springer-Verlag, Berlin, 1971.
%\bibitem[GH94]{GH94} Phillip Griﬃths and Joseph Harris.  {\it Principles of algebraic geometry.} Wiley Classics Library. John Wiley and Sons, Inc., New York, 1994. Reprint of the 1978 original.
%\bibitem[Har77]{Har77} Robin Hartshorne.  {\it Algebraic geometry.} Springer-Verlag, New York-Heidelberg, 1977. Graduate Texts in Mathematics, No. 52.
%\bibitem[Har11]{Har11} Gunter Harder.  {\it Lectures on algebraic geometry II.} Aspects of Mathematics, E39. Vieweg + Teubner, Wiesbaden, 2011. Basic concepts, coherent cohomology, curves and their Jacobians.
%\bibitem[HK00]{HK00} Yi Hu and Sean Keel.  {\it Mori dream spaces and GIT.} Michigan Math. J., 48:331– 348, 2000. Dedicated to William Fulton on the occasion of his 60th birthday.
\bibitem[HM13]{HM13} Christopher D. Hacon, James McKernan. {\it  The Sarkisov program.}
J. Algebraic Geom. 22 (2013), no. 2, 389–405. 
%\bibitem[HMX13]{HMX13} Christopher D. Hacon, James McKernan, and Chenyang Xu.  {\it On the birational automorphisms of varieties of general type.} Ann. of Math. (2), 177(3):1077– 1111, 2013.
\bibitem[HMX14]{HMX14} Christopher D Hacon, James McKernan, and Chenyang Xu.  {\it Boundedness of moduli of varieties of general type.} arXiv preprint arXiv:1412.1186, 2014.
\bibitem[HX13]{HX13} Christopher D. Hacon and Chenyang Xu.  {\it Existence of log canonical closures.} Invent. Math., 192(1):161–-195, 2013.
\bibitem[HX15]{HX15} Christopher D. Hacon and Chenyang Xu.  {\it On the three dimensional minimal model program in positive characteristic.} J. Amer. Math. Soc., 28(3):711-–744, 2015.
\bibitem[Kaw94]{Kaw94} Yujiro Kawamata.  {\it Semistable minimal models of threefolds in positive or mixed characteristic.} J. Algebraic Geom., 3(3):463–-491, 1994.
%\bibitem[Kaw99]{Kaw99} Yujiro Kawamata.  {\it Index 1 covers of log terminal surface singularities.} J. Algebraic Geom., 8(3):519–527, 1999.
%\bibitem[KK94]{KK94} J\'anos Koll\'ar and S\'andor Kov\'acs. Birational geometry of log surfaces. preprint, 1994.
\bibitem[KM98]{KM98} J\'anos Koll\'ar and Shigefumi Mori.  {\it Birational geometry of algebraic varieties,} volume 134 of Cambridge Tracts in Mathematics. Cambridge University Press, Cambridge, 1998. With the collaboration of C. H. Clemens and A. Corti, Translated from the 1998 Japanese original.
%\bibitem[KMM85]{KMM85} Yujiro Kawamata, Katsumi Matsuda, and Kenji Matsuki.  {\it Introduction to the minimal model problem.} In Algebraic geometry, Sendai, 1985, volume 10 of Adv. Stud. Pure Math., pages 283–360. North-Holland, Amsterdam, 1987.
%\bibitem[Kol13]{Kol13} J\'anos Koll\'ar.  {\it Singularities of the minimal model program,} volume 200 of Cambridge Tracts in Mathematics. Cambridge University Press, Cambridge, 2013. With a collaboration of S\'andor Kov\'acs.
\bibitem[KU85]{KU85} Toshiyuki Katsura and Kenji Ueno.  {\it On elliptic surfaces in characteristic $p$.} Math. Ann., 272(3):291-–330, 1985.
\bibitem[La83]{La83} W. Lang, {\it On Enriques surfaces in characteristic $p$. I}, Math. Ann. 265 (1983), no. 1, pp. 45-65.
%\bibitem[Lai11]{Lai11} Ching-Jui Lai. Varieties ﬁbered by good minimal models. Math. Ann., 350(3):533–547, 2011.
%\bibitem[Laz04]{Laz04} Robert Lazarsfeld.  {\it Positivity in algebraic geometry. I,} volume 48 of Ergebnisse der Mathematik und ihrer Grenzgebiete. 3. Folge. A Series of Modern Surveys in Mathematics [Results in Mathematics and Related Areas. 3rd Series. A Series of Modern Surveys in Mathematics]. Springer-Verlag, Berlin, 2004. Classical setting: line bundles and linear series.
%\bibitem[Liu02]{Liu02} Qing Liu.  {\it Algebraic geometry and arithmetic curves,} volume 6 of Oxford Graduate Texts in Mathematics. Oxford University Press, Oxford, 2002. Translated from the French by Reinie Ern´e, Oxford Science Publications.
%\bibitem[Nak04]{Nak04} Noboru Nakayama.  {\it Zariski-decomposition and abundance,} volume 14 of MSJ Memoirs. Mathematical Society of Japan, Tokyo, 2004.
%\bibitem[SC11]{SC11} Vyacheslav Shokurov and Sung Rak Choi.  {\it Geography of log models: theory and applications.} Cent. Eur. J. Math., 9(3):489–534, 2011.

\bibitem[Siu98]{Siu98} Yum-Tong Siu, {\it Invariance of plurigenera.} Invent. Math., 134(3):661-–673, 1998.
\bibitem[Siu02]{Siu02} Yum-Tong Siu. {\it Extension of twisted pluricanonical sections with plurisubharmonic weight and invariance of semipositively twisted plurigenera for manifolds not necessarily of general type.} In Complex geometry (G\"ottingen, 2000), pages 223–-277. Springer, Berlin, 2002.

\bibitem[Suh08]{Suh08} J. Suh, {\it   Plurigenera of general type surfaces in mixed characteristic.} Compos. Math., 144(5):1214–-1226, 2008.
%\bibitem[Suh]{Suh} J. Suh, {\it Deformations and liftings in the minimal model program} Personal communication 7-31-2014.
\bibitem[Tan14] {Tan14}Hiromu Tanaka, {\it Minimal models and abundance for positive characteristic log surfaces.} Nagoya Math. J., 216:1-–70, 2014.
%\bibitem[Tan15a]{Tan15a} Hiromu Tanaka, {\it  Behavior of canonical divisors under purely inseparable base changes.} arXiv preprint arXiv:1502.01381, 2015.
%\bibitem[Tan15b]{Tan15b} Hiromu Tanaka, {\it  Minimal model theory for surfaces over an imperfect ﬁeld.} arXiv preprint arXiv:1502.01383, 2015.
%\bibitem[Tan15c]{Tan15c} Hiromu Tanaka, {\it  The trace map of Frobenius and extending sections for threefolds.} Michigan Math. J., 64(2):227–261, 2015.
\bibitem[Tan15]{Tan15} Hiromu Tanaka, {\it  The X-method for klt surfaces in positive characteristic.} J. Algebraic Geom., 24(4):605–-628, 2015.
\bibitem[Tan16]{Tan16} Hiromu Tanaka, {\it  Minimal model programme for excellent surfaces.} ArXiv:1608.07676
%\bibitem[Ter99]{Ter99} Hiroyuki Terakawa, {\it The d-very ampleness on a projective surface in positive characteristic.} Paciﬁc J. Math., 187(1):187–199, 1999.
%\bibitem[Wal15]{Wal15} Joe Waldron, {\it  Finite generation of the log canonical ring for 3-folds in char $p$.} arXiv preprint arXiv:1503.03831, 2015.
\end{thebibliography}

\documentclass[12pt]{amsart} 
\usepackage{amsfonts,graphics,amsmath,amsthm,amsfonts,amscd, amssymb,amsmath,latexsym}
\usepackage{epsfig}
\usepackage{flafter}

\theoremstyle{plain}
\newtheorem{theorem}{Theorem}[section]
\newtheorem{corollary}[theorem]{Corollary}
\newtheorem{lemma}[theorem]{Lemma}
\newtheorem{claim}[theorem]{Claim}

\newtheorem{proposition}[theorem]{Proposition}
\newtheorem{definition-lemma}[theorem]{Definition-Lemma}

\newtheorem{remark}[theorem]{Remark}

\def\ideal#1.{I_{#1}}
\def\ring#1.{\mathcal {O}_{#1}}
\def\fring#1.{\hat{\mathcal {O}}_{#1}}
\def\proj#1.{\mathbb P(#1)}
\def\pr #1.{\mathbb P^{#1}}
\def\af #1.{\mathbb A^{#1}}
\def\Hz #1.{\mathbb F_{#1}}
\def\Hbz #1.{\overline{\mathbb F}_{#1}}
\def\pic#1.{\operatorname {Pic}\,(#1)}
\def\pico#1.{\operatorname{Pic}^0(#1)}
\def\picg#1.{\operatorname {Pic}^G(#1)}
\def\ner#1.{NS (#1)}
\def\rdown#1.{\llcorner#1\lrcorner}
\def\rup#1.{\ulcorner#1\urcorner}
\def\cone#1.{\operatorname {NE}(#1)}

\def\ccone#1.{\overline{\operatorname {NE}}(#1)}
\def\coef#1.{\frac{(#1-1)}{#1}}
\def\vit#1.{D_{\langle #1 \rangle}}
\def\mm#1.{\overline {M}_{0,#1}}
\def\H1#1.{H^1(#1,{\ring #1.})}
\def\ac#1.{\overline {\mathbb F}_{#1}}

\def\adj#1.{\frac {#1-1}{#1}}
\def\spn#1.{\overline{#1}}
\def\ses#1.#2.#3.{0\to #1\to #2\to #3 \to 0}
\def\pek#1.#2.{\Cal P^{#1}(#2)}
\def\plk#1.#2.{\Cal P^{\leq #1}(#2)}
\def\ev#1.{\operatorname{ev_{#1}}}
\def\bminv#1.{(\nu_1,s_1;\nu_2,s_2;\dots ;\nu_{#1},s_{#1};\nu_{r+1})}
\def\zinv#1.{(\nu_1,s_1;\nu_2,s_2;\dots ;\nu_{#1},s_{#1};0)}
\def\iinv#1.{(\nu_1,s_1;\nu_2,s_2;\dots ;\nu_{#1},s_{#1};\infty)}
\def\map#1.#2.{#1 \longrightarrow #2}
\def\rmap#1.#2.{#1 \dasharrow #2}
\def\emb#1.#2.{#1 \hookrightarrow #2}

\def\N{\mathbb N}

\def\e{\Cal E}

\def\e1{E_1}
\def\e2{E_2}

\def\bB{\mathbf B}

\newcommand\Q{{\mathbb{Q}}}
\newcommand\R{{\mathbb{R}}}
\begin{document}
\title{Invariance of certain plurigenera for surfaces in mixed
 characteristics}
\author{Andrew Egbert}

\author{Christopher D. Hacon} 
\address{Department of Mathematics \\  
University of Utah\\  
Salt Lake City, UT 84112, USA}
 \thanks{The second author
 was supported by NSF research grants no: DMS-1300750, DMS-1265285 and by
  a grant from the Simons Foundation; Award Number: 256202.}

\maketitle
\begin{abstract} We prove the deformation invariance of Kodaira dimension and of certain plurigenera for log surfaces which are smooth over a DVR. \end{abstract}
\section{Introduction}
If $f:X\to T$ is a smooth projective morphism of complex quasi-projective varieties, then by a celebrated theorem of Siu (\cite{Siu98}, \cite{Siu02}), it is known that the plurigenera of the fibers $P_m(X_t):=h^0(mK_{X_t})$ are independent of the point $t\in T$. This result (and its generalizations to log pairs)  is a fundamental fact of great importance in higher dimensional birational geometry. It plays a fundamental role in the construction of moduli spaces of varieties of log general type. Unluckily, this result does not generalize even to families of surfaces over a curve (or a discrete valuation ring DVR).
In \cite{La83}, it is shown that $P_1$ is not deformation invariant for Enriques surfaces in characteristic $2$. In \cite{KU85}, it is shown that in fact the deformation invariance of plurigenera does not hold for certain elliptic surfaces and in \cite{Suh08}, there are examples of smooth families of surfaces of general type over any DVR for which $P_1$ is non constant (and in fact its value can jump by an arbitrarily big amount). On the positive side, in \cite{KU85} it is shown that 
if $X\to {\rm Spec }(R)$ is a smooth family of surfaces over a DVR in positive or mixed characteristic, then one can run the minimal model program for $X$ (over $R$). As a consequence of this, it is observed that $\kappa (X_K)=\kappa (X_k)$ where $k$ is the residue field and $K$ is the fraction field of $R$. It should be noted that the minimal model program is established for semistable families of surfaces in positive or mixed characteristic  (see  \cite{Kaw94}), for log canonical surfaces over excellent base schemes (see \cite{Tan16}) and for $3$ folds over a field $k$ of characteristic $p\geq 7$ (see \cite{HX13} and \cite{Bir15}).
In this paper (Theorem \ref{t-definv} and Corollary \ref{c-mmp}), we generalize the result of Katsura and Ueno to log surfaces (smooth over a DVR) and we show the deformation invariance of certain plurigenera.
\begin{theorem}Let $(X,B)$ be a klt pair which is log smooth, projective of dimension $2$ over a DVR $R$ with perfect residue field $k$ of characteristic $p > 0$, then $\kappa (K_{X_k}+B_k)=\kappa (K_{X_K}+B_K)$  and we may run the minimal model program with scaling of an ample divisor $H$. 
Moreover, if either $\kappa (K_{X_k}+B_k)\ne 1$ or $\kappa (K_{X_k}+B_k)= 1$ and $B_k$ is big over ${\rm Proj}R(K_{X_k}+B_k)$, then there exists an integer $m_0>0$  such that for any positive integer $m\in m_0\N$ we have
$$h^0(m (K_{X_K}+B_K))=h^0(m (K_{X_k}+B_k)).$$\end{theorem}
The strategy is to reduce the proof of the above theorem to the case when $(X_k,B_k)$ is terminal and $\mathbf B (K_{X_k}+B_k)$ contains no components of the support of $B_k$. In this case we observe that the steps of a $K_{X_k}+B_k$ mmp are also steps of a $K_{X_k}$ mmp and we are thus able to deduce the result from \cite{KU85}.
\begin{remark} Many results and techniques in this paper were developed in the first author's Ph.D.'s thesis \cite{Egb16}.\end{remark}
%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\section{Preliminaries}
Let $X$ be a normal projective variety, ${\rm WDiv} (X)$ the group of Weil divisors. If $B=\sum b_iB_i\in {\rm WDiv} _\Q(X)$ is a $\Q$-divisor on $X$, then $\lfloor B \rfloor=\sum \lfloor b_i\rfloor B_i$
where $\lfloor b_i\rfloor={\rm max}\{n\in \mathbb Z|n\leq b_i\}$. We denote $\{B\}=B-\lfloor B \rfloor$ and $|B|=|\lfloor B \rfloor|+\{B\}$ where $$|\lfloor B \rfloor|=\{D\in {\rm WDiv} (X)|
 D\geq 0,\ D-\lfloor B \rfloor=(f),\ f\in K(X)\}.$$ The {\bf stable base locus} of $B$ is $\bB (D)=\cap _{m\in \N} {\rm Bs}(mD)$.
If $D_1,\ldots , D_r$ are $\Q$-divisors on a normal projective variety $X$, then $$R(D_1,\ldots , D_r):=\bigoplus _{m_1,\ldots , m _r\in \N}H^0(\mathcal O _X(\sum _{i=1}^r m_iD_i)).$$ 
Let $(X,B)$ be a {\bf pair} so that $X$ is normal, $0\leq B$ is a $\Q$-divisor and $K_X+B$ is $\Q$-Cartier. If $\nu :X'\to X$ is a proper birational morphism, then we write $K_{X'}+B_{X'}=\nu ^*(K_X+B)$. We say that $(X,B)$ is {\bf Kawamata log terminal or klt} (res. {\bf terminal}) if for any log resolution $\lfloor B_{X'}\rfloor \leq 0$ (resp. $\lfloor B_{X'}\rfloor \leq E$ where $E$ denotes the reduced exceptional divisor). 
We let $\mathbf M _B$ be the $b$-divisor defined by the sum of the strict transform of $B$ and the exceptional divisors (over $X$). 
We refer the reader to \cite{KM98} and \cite{BCHM10} for the standard definitions of the minimal model program including extremal rays, flipping and divisorial contractions, running a minimal model program with scaling, log terminal and weak log canonical models.

\begin{theorem}\label{t-ct} Let $(X,B)$ a 2-dimensional projective klt pair over an algebraically closed field $k$. Then 
$$\overline {NE}(X)=\overline {NE}(X)_{K_X+B\geq 0}+\sum _{i\in I}\R _{\geq 0}C_i$$
where $I$ is countable, $(K_X+B)\cdot C_i<0$, $C_i$ is rational and $C_i^2<0$.
If $H$ is an ample $\Q$-divisor on $X$, then the set $\{ i\in I|(K_X+B+H)\cdot C_i\leq 0\}$ is finite.
\end{theorem}
\begin{proof} See \cite[3.13, 3.15]{Tan14}.
\end{proof}
\begin{lemma}\label{l-dc} Let  $X$ be a surface over an algebraically closed field $k$ and $(X,B)$ a projective klt pair.
If $R$ is a $K_X+B$ negative extremal ray, then there exists a proper morphism $f:X\to X'$ such that $f_*\mathcal O _X=\mathcal O _{X'}$, $f$ contracts a curve $C\subset X$ if and only if $[C]=R$.\end{lemma}
\begin{proof} See \cite[3.21]{Tan14}.
\end{proof}
\begin{theorem}\label{t-mmps} Let $X$ be a projective surface over a perfect field $k$. Assume that $(X,B)$ is klt and $H$ is an ample $\mathbb Q$-divisor on $X$. Then \begin{enumerate}
\item the ring $R(K_X+B,K_X+B+H)$ is finitely generated, and
\item if $K_X+B$ is pseudo-effective, then there exists a birational morphism $\nu :X\to X'$ and a constant $\epsilon >0$ such that $\nu$ is a $K_X+B+tH$ minimal model for any $0\leq t\leq \epsilon$.\end{enumerate}\end{theorem}
\begin{proof} By \cite{Tan14}, the $K_X+B$ mmp with scaling of $H$ terminates. Therefore there is a sequence of rational numbers $t_0\leq t_1\leq \ldots \leq t_n$ and morphisms $f_i:X\to X_i$ (induced by successive divisorial contractions) such that \begin{enumerate} \item if $t_0>0$, then $\kappa (K_X+B+tH)<0$ for any $0\leq t< t_0$,
\item $R(K_X+B+t_{0}H,
K_X+B+t_iH)\cong R( K_{X_i}+B_i+t_{0}H_i,
K_{X_i}+B_i+t_iH_i)$ where $B_i=f_{i,*}B$ and $H_i=f_{i,*}H$,
\item  $K_{X_i}+B_i+t_{i-1}H_i$ and
$K_{X_i}+B_i+t_iH_i$ are both nef (and hence semiample).\end{enumerate} 
It then follows easily that $$R:= \oplus _{i=1}^n R(K_{X_i}+B_i+t_{i-1}H_i,K_{X_i}+B_i+t_iH_i)$$ is finitely generated (cf. \cite[2.7]{HX15}).
Since $R$ surjects on to $R(K_X+B,K_X+B+H)$, this ring is also finitely generated. 

Finally, if $K_X+B$ is pseudo-effective, then $t_0=0$, $K_{X_1}+B_1+tH_1$ is nef for $0\leq t\leq t_1$ and $X\to X_1$ is given by a sequence of $K_X+B+tH$ divisorial contractions for any $0\leq t <t_1$.\end{proof}  

\begin{proposition}\label{p-mmps} Let $X$ be a projective surface over an algebraically closed field $k$. Assume that $(X,B)$ is a klt pair and $\nu :X'\to X$ is a proper birational morphism such that $(X',B')$ is terminal where $K_{X'}+B'=\nu ^*(K_X+B)$. Let $\Theta =B'-B'\wedge N_\sigma (K_{X'}+B')$ and $\phi ':X'\to X'_M$ the minimal model for $(X',\Theta )$. If $\phi:X\to X_M$ is the minimal model for $(X,B)$, then the rational map $\mu:X'_M\to X_M$ is a morphism and $K_{X'_M}+\phi '_*\Theta =\mu ^*(K_{X_M}+\phi _* B)$.
If $\kappa (K_X+B)=1$ and $B$ is big over ${\rm Proj}R(K_X+B)$, then $\Theta $ is big over ${\rm Proj}R(K_{X'}+\Theta)$.
\end{proposition}
\begin{proof} Consider the morphism $\psi:X'\to X_M$. Since $K_X+B=\phi ^*(K_{X_M}+\phi _*B)+E$, then $K_{X'}+B'=\psi ^*(K_{X_M}+\phi _*B) +\nu ^*E$ where $K_{X_M}+\phi _*B$ is nef and $\nu ^*E$ is effective and $\psi$ exceptional. It follows that 
$N_\sigma (K_{X'}+B')=\nu ^*E$ and so $K_{X'}+\Theta =\psi ^* (K_{X_M}+\phi _*B)+E'$ where $0\leq E'\leq \nu ^*E$. In particular the divisors contracted by $\phi '$ are precisely the divisors contained in ${\rm Supp} (E')$ and so $X'\to X_M$ factors through $\phi '$. 
We have $K_{X'_M}+\phi '_* \Theta =\mu ^*(K_{X_M}+\phi _*B)+\phi '_* E'$ where $\mu _* (\phi '_*E')\leq \phi _* E=0$ and hence $\phi'_*E $ is $\mu$ exceptional. By the negativity lemma, it follows that $\phi '_* E'=0$.

Note that since $H^0(m(K_X+B))\cong H^0(m(K_{X'}+\Theta ))$ for all $m\geq 0$, it follows that $Z:={\rm Proj}R(K_X+B)={\rm Proj}R(K_{X'}+\Theta)$. The bigness of $B$ over $Z$ is equivalent to $B\cdot X_z>0$ for general $z\in Z$. But then
$$\Theta \cdot X'_z=\mu _* \phi '_*\Theta \cdot (X'_M)_z=\phi _*B\cdot (X'_M)_z=B\cdot X_z>0$$ and so $\Theta$ is big over $Z$.
\end{proof}

In what follows $R$ will denote a DVR with residue field $k$ and fraction field $K$. If $f:X\to {\rm Spec}(R)$ is a morphism, then we let $X_K=X\times _{{\rm Spec}(R)}{\rm Spec}(K)$ be the generic fiber and $X_k= X\times _{{\rm Spec}(R)}{\rm Spec}(k)$ be the special fiber. We say that a pair $(X,B)$ is log smooth over $R$ if $X$ and each strata of the support of $B$ are smooth over ${\rm Spec}( R)$.
%\begin{lemma}\label{l-cbc} Let $f:X\to {\rm Spec }(R)$ be a projective morphism from a normal variety to a DVR. If $\FF$ is a coherent sheaf on $X$ which is flat over ${\rm Spec}(R)$, then $\chi _k(\FF _k)=\chi _K(\FF _K)$. Moreover, for any $i\geq 0$ the following are equivalent. \begin{enumerate}\item $\dim _kH^i(\FF _k)=\dim _KH^i(\FF _K)$, \item $H^i(\FF )$ is free over ${\rm Spec}(R)$ and $H^i(\FF )\otimes _R k\to H^i(\FF _k)$ is an isomorphism, \item $H^{i+1}(\FF)$ is a torsion free $R$-module. \end{enumerate} \end{lemma} \begin{proof} \cite[5.3.20, 5.3.22]{Liu02}. [xxxcheck] \end{proof}
\begin{lemma}\label{l-open} Let $f:X\to {\rm Spec}(R)$ be a smooth projective morphism from a normal variety to a DVR and $L$ a line bundle on $X$, then $L$ is ample (resp. nef) if and only if so is $L_k:=L|_{X_k}$.\end{lemma}
\begin{proof} Clearly, if $L$ is ample or nef, then so is $L_k$. It is well known that ampleness is an open condition and so if $L_k$ is ample then so is $L$. Finally if $L_k$ is nef and $H$ is ample, then $L_k+tH_k$ is ample for any $t>0$, so that $L+tH$ is ample and hence $L$ is nef.
\end{proof}
\begin{lemma} Let $(X,B)$ be a log pair which is log smooth over ${\rm Spec }(R)$ where $R$ is a DVR. If $R\subset \tilde R$ is an inclusion of DVR's, then $(X_{\tilde R},B_{\tilde R})$ is log smooth over $ {\rm Spec }(\tilde  R)$. If $(X,B)$ is terminal (resp. klt), then so is $(X_{\tilde R},B_{\tilde R})$.\end{lemma}
\begin{proof}Since smoothness is preserved by base change, it follows that $(X_{\tilde R},B_{\tilde R})$ is log smooth over $ {\rm Spec }(\tilde R)$. The pair $(X,B)$ is klt (resp. terminal) if and only if the coefficients of $B$ are $<1$ (resp.  the coefficients of $B$ are $<1$ and if two components intersect, then the sum of the coefficients is $<1$). Since there is a one to one correspondence of components of $B$ with components of $B_k$, the lemma follows. \end{proof}
\begin{theorem}\label{t-KU} [Katsura-Ueno \cite{KU85}] Let $f:X\to {\rm Spec}(R)$ be an algebraic space which is proper, separated of finite type and two dimensional over $ {\rm Spec}(R)$ where $R$ is a DVR with algebraically closed residue field $k$ and field of fractions $K$. If $X_k$ contains an exceptional curve of the first kind $e\subset X_k$ then there exists a DVR $\tilde R\supset R$ with residue field $k$ and a surjective proper morphism $\pi :X\times {\rm Spec}(\tilde R)\to \tilde Y$ over ${\rm Spec}(\tilde R)$ where $\tilde Y\to {\rm Spec}(\tilde R)$  is proper, separated of finite type and two dimensional, $\pi _k$ contracts the exeptional curve of the first kind $e\subset X_k$ and $\pi _K:X_{\tilde K}\to \tilde Y_{\tilde K}$ is also a contraction of an exeptional curve of the first kind.\end{theorem}

%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%%
\section{Main result}
\begin{theorem}\label{t-definv} Let $(X,B)$ be a klt pair which is log smooth, projective of dimension $2$ over a DVR $R$ with perfect residue field $k$ of characteristic $p>0$ and perfect fraction field $K$. If $K_X+B$ is $\Q$-Cartier then  $\kappa (K_{X_k}+B_k)=\kappa (K_{X_K}+B_K)$ and if either $\kappa (K_{X_k}+B_k)\ne 1$ or $\kappa (K_{X_k}+B_k)= 1$ and $B_k$ is big over ${\rm Proj}R(K_{X_k}+B_k)$, then there exists an integer $m_0$ such that for any positive integer  $m\in m_0\N$ we have
$$h^0(m (K_{X_K}+B_K))=h^0(m (K_{X_k}+B_k)).$$
\end{theorem}
\begin{proof} Consider an inclusion of DVR's $R\subset \tilde R$. If $\tilde k$ and $\tilde K$ denote the residue field and the fraction field of $\tilde R$, then $h^0(m (K_{X_k}+B_k))=h^0(m (K_{ X_{\tilde k}}+B_{\tilde k}))$ and $h^0(m (K_{X_K}+B_K))=h^0(m (K_{ X_{\tilde K}}+B_{\tilde K}))$. Thus we are free to replace $X\to R$ by a base  extension $\tilde X=X\times _R\tilde R\to \tilde R$. In particular we may assume that $k$ is algebraically closed.

If $h^0(m (K_{X_k}+B_k))=0$, then by semicontinuity $h^0(m (K_{X_K}+B_K))=0$. Therefore, the theorem holds trivially in the case $\kappa (K_{X_k}+B_k)=-\infty $.  Thus we may assume that $\kappa (K_{X_k}+B_k)\geq 0$.

\begin{claim}\label{c-1} The theorem holds under the additional assumption that $(X_k,B_k)$ is terminal and no component of the support of $B_k$ is contained in $\bB (K_{X_k}+B_k)$.\end{claim}
\begin{proof} Since $k$ is algebraically closed, then by the Cone Theorem (Theorem \ref{t-ct}) $$\overline {NE}(X_k)=\overline {NE}(X_k)_{K_{X_k}+B_k\geq 0}+\sum _{i\in I}\R _{\geq 0}C_i$$
where $I$ is countable, $(K_{X_k}+B_k)\cdot C_i<0$ and $C_i$ is rational and not  contained in the support of $B_k$. 
Notice in fact that if $C_i$ is contained in the support of $B_k$, then since  $(K_{X_k}+B_k)\cdot C_i<0$ we have $C_i\subset \bB (K_{X_k}+B_k)$, which we have assumed is impossible. Note then that $C_i\cdot B_k\geq 0$ and so $K_{X_k}\cdot C_i<0$ 
and hence if $C_i$ spans a $K_{X_k}+B_k$-negative extremal ray, then it also spans a  $K_{X_k}$-negative extremal ray and so it  can be contracted by a divisorial contraction of an exceptional curve of the first kind $X_k\to X'_k$. In particular $X'_k$ is 
also a smooth surface. %(Lemma \ref{l-dc}).  But then, since $X_k$ is terminal, so is $X'_k$ and in particular $X'_k$ is smooth. 
Thus we may assume that $C_i$ is an extremal curve of the first kind. By Theorem \ref{t-KU} (after extending $R$), we may assume that there is a morphism $X\to X'$ 
of smooth surfaces over $R$ such that $X_K\to X'_K$ also contracts  an extremal curve of the first kind.

Suppose now that $\nu :X\to \bar X$ is a morphism of smooth surfaces over ${\rm Spec }(R)$ such that $X_K\to \bar X_K$ and $X_k\to \bar X_k$ are given by a finite sequence of contractions of  extremal curves of the first kind such that the exceptional locus of $X_k\to \bar X_k$ contains no components of $B_k$.
Then $(\bar X_k,\bar B_k)$ is terminal and $K_{X_k}+B_k=\nu _k^*(K_{\bar X_k,}+\bar B_k)+F_k$ where $B_k=\nu _{k,*}^{-1}\bar B_k$ and $B_k\wedge F_k=0$. In particular $\bB (K_{ X_k}+ B_k)=\bB (K_{\bar X_k}+\bar B_k)+F_k$.
 Suppose that $C\subset \bar X_k$
is contained in $\bB (K_{\bar X_k}+\bar B_k)\cap {\rm Supp}(\bar B_k)$, then $\nu ^{-1}_*C\subset \bB (K_{X_k}+B_k)\cap {\rm Supp}(B_k)$ which is impossible. Therefore, if $K_{\bar X_k}+\bar B_k$ is not nef, we can continue to contract exceptional curves of the first kind. Since each contraction reduces the Picard number of the central fiber $X_k$ by one, this procedure must terminate after finitely many steps. We may therefore assume that $K_{\bar X_k}+\bar B_k$ is semiample. In particular $K_{\bar X_k}+\bar B_k$ is nef and hence so is $K_{\bar X}+\bar B$ (see Lemma \ref{l-open}).

Suppose now that $\nu (K_{X_k}+B_k)=2$. In this case $K_{\bar X_k}+\bar B_k$ is nef and big so that there exists an integer $m_0>0$ such that $h^i(m(K_{\bar X_k}+\bar B_k))=0$ for all $m\in m_0\N$ and all $i>0$ (see \cite[2.6]{Tan15}). By semicontinuity, we also have $h^i(m(K_{\bar X_K}+\bar B_K))=0$ for all $m\in m_0\N$. But then, by flatness,
$$h^0(m(K_{X_K}+ B_K))=h^0(m(K_{\bar X_K}+\bar B_K))=\chi (m(K_{\bar X_K}+\bar B_K))=\qquad$$ $$\qquad \chi (m(K_{\bar X_k}+\bar B_k))=h^0(m(K_{\bar X_k}+\bar B_k))=h^0(m(K_{X_k}+ B_k)).$$

Suppose that $\kappa (K_{X_k}+B_k)=0$. Then we have $K_{\bar X_k}+\bar B_k\sim _\Q 0$. By Lemma \ref{l-open}, it follows that $\pm (K_{\bar X_K}+\bar B_K)$ is nef  and hence that $K_{\bar X_K}+\bar B_K\equiv 0$. By \cite[1.2]{Tan14}, $K_{\bar X_K}+\bar B_K\sim _\Q 0$. Thus there exists an integer $m_0>0$ such that $m_0(K_{\bar X_K}+\bar B_K)\sim 0$ and $m_0(K_{\bar X_k}+\bar B_k)\sim 0$.
Thus $h^0(m(K_{X_K}+ B_K))=h^0(m(K_{X_k}+ B_k))$  for all $m\geq 0$ divisible by $m_0$.

Suppose that $\kappa (K_{X_k}+B_k)=1$. Since $ K_{\bar X_k}+\bar B_k$ is nef, so is $K_{\bar X_K}+\bar B_K$. In particular $\kappa (K_{\bar X_K}+\bar B_K)\geq 0$ and thus, by semicontinuity we have  $\kappa (K_{\bar X_K}+\bar B_K)\in \{0,1\}$. 
 Let $H$ be a sufficiently ample divisor on $\bar X$. Then 
$(K_{\bar X_K}+\bar B_K)\cdot H_K=(K_{\bar X_k}+\bar B_k)\cdot H_k>0$ so $K_{\bar X_K}+\bar B_K\not\equiv 0$.
Therefore  $\kappa (K_{\bar X_K}+\bar B_K)= 1$.

 Finally, suppose that  $\kappa (K_{X_k}+B_k)=1$ and $B_k$ is big over ${\rm Proj}R(K_{X_k}+B_k)$. 
Note that  $\bar B_k$ is also big over ${\rm Proj}R(K_{X_k}+B_k)$ and hence $\bar B_k+K_{\bar X_k}+\bar B_k$ is big. Thus
we may write $\bar B_k+K_{\bar X_k}+\bar B_k\sim _\Q \bar A_k+\bar E_k$ where $\bar A_k$ is ample and $\bar E_k$ is effective. For any rational number $0<\epsilon \ll 1$, the pair $(\bar X _k, \Delta _k=(1-\epsilon)\bar B_k+\epsilon \bar E_k)$ is Kawamata log terminal and so the corresponding multiplier ideal sheaf is trivial
$\mathcal J (\Delta_k)=\mathcal O _{\bar X _k}$. If $L=N=m(K_{\bar X_k}+\bar B_k)$, then $N$ is nef and not numerically equivalent to zero while $L-( K_{\bar X_k}+\Delta_k)\sim _\Q (m-1-\epsilon )(K_{\bar X_k}+\bar B_k)+  \epsilon \bar A_k$ is ample and so by \cite[0.3]{Tan15},
$H^i(\mathcal O _{\bar X_k}(m(l+1)(K_{\bar X_k}+\bar B_k)))=0$ for $i>0$ and $l\gg 0$.  By semicontinuity, $H^i(\mathcal O _{\bar X_K}(m(l+1)(K_{\bar X_K}+\bar B_K)))=0$ for $i>0$ and $l\gg 0$ and hence $h^0(\mathcal O _{\bar X_k}(m(l+1)(K_{\bar X_k}+\bar B_k)))=h^0(\mathcal O _{\bar X_K}(m(l+1)(K_{\bar X_K}+\bar B_K)))$.
\end{proof}
We will now consider the general case. 
Since $(X,B)$ is log smooth over $R$, there is a sequence of blow ups along strata of $\mathbf M_B$ say $\nu : X'\to X$ such that $K_{X'}+B'=\nu ^*(K_X+B)$ is terminal and in particular $B'\geq 0$ and $(X',B')$ is log smooth. Since $R(K_{X'_k}+B'_k)\cong R(K_{X_k}+B_k)$ is finitely generated, $N_\sigma (K_{X'_k}+B'_k)$ is a $\Q$-divisor and hence so is $$\Theta _k:=B'_k-(B'_k\wedge N_\sigma (K_{X'_k}+B'_k)).$$
Note that $R(K_{X'_k}+\Theta _k)\cong R(K_{X'_k}+B'_k)$, $(X'_k, \Theta _k)$ is terminal and no component of $\Theta _k$ is contained in $\bB (K_{X'_k}+\Theta _k)$ \cite[2.8.3]{HMX14} and \cite[2.4]{HX13}.
Let $\Theta $ be the unique $\Q$-divisor supported on $B'$ such that $\Theta |_{X'_k}=\Theta _k$.
We remark that if $\kappa (K_{X_k}+B_k)=1$ and $B_k$ is big over ${\rm Proj}R(K_{X_k}+B_k)$, then by Proposition \ref{p-mmps}, $\Theta _k$ is big over  ${\rm Proj}R(K_{X'_k}+\Theta _k)$.
By Claim \ref{c-1}, it follows that $\kappa (K_{X'_K}+\Theta _K)=\kappa  (K_{X'_k}+\Theta _k)$ and 
there exists an integer $m_0>0$ such that  $$h^0(m(K_{X'_K}+\Theta _K))=h^0(m(K_{X'_k}+\Theta _k))\qquad \forall m\in m_0\mathbb N.$$ 
By semicontinuity, we then have $$h^0(m(K_{X_k}+B_k))\geq h^0(m(K_{X_K}+B_K))\geq h^0(m(K_{X'_K}+B'_K))\geq\qquad$$ $$\qquad h^0(m(K_{X'_K}+\Theta _K))=h^0(m(K_{X'_k}+\Theta _k))=h^0(m(K_{X_k}+B_k))$$ and hence
$h^0(m(K_{X_k}+B_k))=h^0(m(K_{X_K}+B_K))$. 
The equality
 $\kappa (K_{X_k}+B_k)=\kappa (K_{X_K}+B_K)$ follows similarly.

\end{proof}
\begin{corollary}\label{c-fingen} Let $(X,B)$ be a klt pair which is log smooth, projective of dimension $2$ over a DVR $R$ with perfect residue field $k$ of characteristic $p>0$ and perfect fraction field $K$. If $K_X+B$ is $\Q$-Cartier, $H$ is an ample divisor on $X$, and either $\kappa (K_{X_k}+B_k)\in \{0,2\}$ or $\kappa (K_{X_k}+B_k)=1$ and $B_k$ is big over ${\rm Proj} R (K_{X_k}+B_k)$, then
$R(K_X+B,K_X+B+H)$ is finitely generated.\end{corollary}
\begin{proof} By Theorem \ref{t-mmps}, $R(K_{X_k}+B_k,K_{X_k}+B_k+H_k)$ is finitely generated and hence there is an sequence of rational numbers $0=q_0<q_1<q_2<\ldots <q_n=1$ and an integer $m$ such that $R(m(K_{X_k}+B_k+q_iH_k),m(K_{X_k}+B_k+q_{i+1}H_k))$ is generated in degree 1 i.e. by $H^0(m(K_{X_k}+B_k+q_iH_k))$ and $H^0(m(K_{X_k}+B_k+q_{i+1}H_k))$.
We may assume that each $m(K_{X}+B+q_iH)$ is Cartier.
By Theorem 
  \ref{t-definv}, after replacing $m$ by a multiple, we may assume that $$H^0(m(K_X+B+q_iH))\to H^0(m(K_{X_k}+B_k+q_iH_k))$$ is surjective. 
Therefore, the induced map $$S^kH^0(m(K_X+B+q_iH))\otimes S^jH^0(m(K_X+B+q_{i+1}H))\to$$ $$S^kH^0(m(K_{X_k}+B_k+q_iH_k))\otimes S^jH^0(m(K_{X_k}+B_k+q_{i+1}H_k))\to$$ $$H^0(mk(K_{X_k}+B_k+q_iH_k)+mj(K_{X_k}+B_k+q_{i+1}H_k))$$ is surjective.
By Nakayama's lemma, $$S^kH^0(m(K_{X}+B+q_iH))\otimes S^jH^0(m(K_{X}+B+q_{i+1}H))\to $$
$$H^0(mk(K_{X}+B+q_iH)+mj(K_{X}+B+q_{i+1}H))$$ is surjective and so $R(m(K_X+B+q_iH),m(K_X+B+q_{i+1}H))$ is finitely generated.  It follows easily that $R(m(K_X+B),m(K_X+B+H))$ is finitely generated. By \cite[2.25]{CL13}  $R(K_X+B,K_X+B+H)$ is finitely generated. 
\end{proof}
\begin{remark}\label{r_fg} By the same arguments in the proof of Corollary \ref{c-fingen}, one sees that if $\kappa (K_{X_k}+B_k)=2$ and $H_1,\ldots ,H_l$ are ample divisors, then $$R(K_X+B, K_X+B+H_1,\ldots , K_X+B+H_l):=$$
$$\bigoplus _{(m_0,\ldots ,m_l)\in \mathbb N^{l+1}}H^0(m_0(K_X+B)+m_1( K_X+B+H_1)+\ldots +m_l(K_X+B+H_l))$$
is finitely generated.
\end{remark}
\begin{corollary}\label{c-mmp} Let $(X,B)$ be a klt pair which is log smooth, projective of dimension $2$ over a DVR $R$ with perfect residue field $k$ of characteristic $p>0$ and perfect fraction field $K$. If $K_X+B$ is $\Q$-Cartier and $H$ is a sufficiently ample divisor on $X$ such that $H$ is general in $N^1(X/R)$,  then (after possibly extending $R$) we may run the $K_X+B$ mmp over $R$ with scaling of $H$ which  terminates with a $K_X+B+\tau H$ minimal model $X\dasharrow \bar X$ over $R$ where $\tau ={\rm inf}\{t\geq 0|\kappa (K_{X_k}+B_k+tH_k)\geq 0\}$.\end{corollary}
\begin{proof} Suppose that $X\dasharrow X_1 \dasharrow \ldots \dasharrow X_i$ is a sequence of steps in the $K_X+B$ mmp over $R$ with scaling of $H$, in particular $X_i$ is $\Q$-factorial and there exists a rational number $\lambda_i \in \Q_{\geq 0}$ such that $K_{X_i}+B_i+\lambda _i H_i$ is nef over $R$ where $B_i$ and $H_i$ are the strict transforms of $B$ and $H$ on $X_i$. 
Let $\lambda _{i+1}={\rm inf}\{\lambda \geq 0|K_{X_i}+B_i+\lambda  H_i\ {\rm is\ nef}\}$. If $\lambda _{i+1}=0$ we are done, and therefore we may assume that $0<\lambda _{i+1}\leq \lambda _i$.

Suppose that $K_{X_k}+B_k+\lambda _{i+1} H_k$ is big. Since bigness is an open condition, $K_{X_k}+B_k+\lambda ' H_k$ is also big for some rational number $\lambda '<\lambda _{i+1} $. By Corollary \ref{c-fingen}, $R(K_X+B+\lambda ' H, K_X+B+\lambda _i H)$ is finitely generated and hence so is $R(K_{X_i}+B_i+\lambda ' H_i, K_{X_i}+B_i+\lambda _{i}H_i)$. It then follows easily that $\lambda _{i+1}\in \Q$.
 Since $R(K_{X_i}+B_i+\lambda _{i+1}H_i)$ is finitely generated, there exists a morphism 
$\mu _i:X_i\to Z_i:={\rm Proj}(R(K_{X_i}+B_i+\lambda _{i+1}H_i))$. 

We claim that $\rho (X_i/Z_i)=1$. To see this, pick ample $\Q$-divisors $H^1,\ldots , H^l$ on $X$ such that $||H-H^i||\ll 1$ in $N^1(X)$ and  $H^1,\ldots , H^l$ span $N^1(X)$.
Let $\mathcal C$ be the cone spanned by $K_X+B+\lambda ' H, K_X+B+H^1,\ldots , K_X+B+H^l$, then by Remark
\ref{r_fg}, the adjoint ring
$$R( K_X+B+\lambda ' H, K_X+B+H^1,\ldots , K_X+B+H^l),$$ is finitely generated.
Let $\mathcal C=\cup _{j=1}^s\mathcal C_j$ be the finite decomposition of $\mathcal C$ in to rational cones given by \cite[3.5]{CL13}. 
Let $K_X+B+\mathcal H$  be  an element in the interior of one of the maximal dimensional cones $\mathcal C _j$, then following \cite[3.3]{HM13} there is an induced birational map $g:X\dasharrow Y={\rm Proj} R(K_X+B+\mathcal H)$ to a log terminal model of $K_X+B+\mathcal H$. Since $H$ is general in $N^1(X)$, it follows that $\lambda _{i+1}$ is determined by the intersection of the segment $[K_X+B+\lambda 'H,K_X+B+H]$ with a codimension one face of the decomposition $\mathcal C=\cup _{j=1}^s\mathcal C_j$. We may assume that this face is shared by maximal dimensional cones $\mathcal C _i$ and $\mathcal C _{i+1}$ where $K_X+B+\lambda _{i+1}H$ corresponds to a general element of $ \mathcal C _{i}\cap \mathcal C _{i+1}$ and  if $K_X+B+\mathcal H\in \mathcal C _i^\circ$, then $X_i\cong {\rm Proj}   R(K_X+B+\mathcal H)$.
Following \cite[3.3(4)]{HM13}, there is a morphism $\mu _i:X_i\to Z_i:={\rm Proj}(R(K_{X_i}+B_i+\lambda _{i+1}H_i))$ of relative Picard number 1.

If ${\rm Ex}(\mu _i)$ is a divisor, then $Z_i$ is $\Q$-factorial, we let $X_{i+1}=Z_i$ and 
$X_i\to X_{i+1}$ is the required divisorial contraction. 
If ${\rm Ex}(\mu _i)$ is not a divisor, then $X_i\to Z_i$ is a flipping contraction and we let $X_{i+1}={\rm Proj}(R(K_{X_i}+B_i+(\lambda _{i+1}-\delta)H_i))$ for any $0<\delta \ll 1$. Note that by the finite generation of $R(K_X+B+\lambda ' H, K_X+B+H)$, it follows that $X_{i+1}$ does not depend on $\delta$. 
After replacing $X_i$ by $X_{i+1}$, we may repeat the above procedure. 

	Suppose now that $K_{X_k}+B_k+\lambda _{i+1} H_k$ is not big and $K_{X_k}+B_k+\lambda _{i} H_k$ is  big. Note that by what we have observed above $X\dasharrow X_{i}$ is a minimal model for $(X,B+\lambda H)$ for any $\lambda _{i+1}<\lambda <\lambda _i$ and so it is a $\Q$-factorial weak log canonical model for $(X,B+\lambda _{i+1} H)$. By construction, $X\dasharrow X_{i}$ is $K_X+B+\lambda _{i+1} H$ negative and hence  $X\dasharrow X_{i}$ is a minimal model for $(X,B+\lambda _{i=1}H)$.
\end{proof}

\end{document}